OpenStudy (nightmarenight):

two questions!! right answer will give medal and fan Given the function T(z) = 7 – z, find T(–4). The number of calories burned, C, varies directly with the time spent exercising, t. When Dennis walks for 4 hours, he burns 800 calories. Show the direct linear variation in an equation

OpenStudy (cggurumanjunath):

T ( Z ) = 7 -Z T (-4) =7- (-4) =7 +4

OpenStudy (nightmarenight):

So it's 11 @CGGURUMANJUNATH

OpenStudy (cggurumanjunath):

:)